The sum of the ages of John and Julia are 107 years. 6 years ago, Johns age was 4 times Julia's age. How old is John NOW???
Answer by addingup(3677)

x+y = 107
y = 107-x
x-6 = 4(y-6)
x-6 = 4y-24
x-6 = 4(107-x)-24
x-6 = 428-4x-24
x-6 = 404-4x
5x = 410
x = 82 This is John's age today. And
Julia: 107-82 = 25
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check - "6 years ago John's age was 4 times Julia's:
82-6 = 76 This is John 6 years ago
25-6 = 19 This is Julia 6 years ago
19*4 = 76 We have the correct answer